the sum of julies age and her fathers age is 58. if julies father is 10 years younger than triple Julie's age, find their age.
Answer by Maths68(1474)

Let
Present age of Julies = j years old
Present age of Father = f years old

the sum of julies age and her fathers age is 58
j+f=58
j=58-f.............(1)
julies father is 10 years younger than triple Julie's age
3j=f-10............(2)
Put the value of j from (1) to (2)
3j=f-10
3(58-f)=f-10
174-3f=f-10
-3f-f=-10-174
-4f=-164
-4f/-4=-164/-4
f=41
Put the value of f in (1)
j=58-f.............(1)
j=58-41
j=17



Present age of Julies = j = 17 years old
Present age of Father = f = 41 years old